Required information [The following information applies to the questions displayed below.) Beech Corporation is a merchandisi1. Estimated sales for July, August, September, and October will be $240,000, $260,000, $250,000, and $270,000, respectively.
Answer #1

1.

Cash Collections budget
July August September Total
Collected in same month (35%) $          84,000 $           91,000 $           87,500 $         262,500
Collected in following month (65%) $       133,000 $         156,000 $         169,000 $         458,000
Total Cash Collections $       217,000 $         247,000 $         256,500 $         720,500

2a.

Merchandise Purchase Budget
July August September Total
Sales $       240,000 $         260,000 $         250,000 $         750,000
Cost of Goods Sold @60% $       144,000 $         156,000 $         150,000 $         450,000
Add : Desired Ending Inventory $          39,000 $           37,500 $           40,500 $           40,500
Total Needs $       183,000 $         193,500 $         190,500 $         490,500
Less Beginning Inventory $          36,000 $           39,000 $           37,500 $           36,000
Merchandise Purchase $       147,000 $         154,500 $         153,000 $         454,500

2b.

Cash Payments budget
July August September Total
Paid in same month (40%) $          58,800 $           61,800 $           61,200 $         181,800
Paid in following month (60%) $          74,000 $           88,200 $           92,700 $         254,900
Total Cash Payments $       132,800 $         150,000 $         153,900 $         436,700

3.

Income Statement
Sales Revenue $       750,000
Cost of Goods Sold $       450,000
Gross Profit $       300,000
Selling and administrative expenses $       108,000
Depreciation $          24,000
Net Income $       168,000

4.

Balance Sheet
Assets
Cash $       266,800 =91000+720500-436700-36000*3
Accounts Receivable $       162,500 =250000*65%
Inventory $          40,500
Plant and Equipment, net $       189,000 =213000-8000*3
Total Assets $       658,800
Liabilities
Accounts Payables $          91,800 =153000*60%
Common Stock $       328,000
Retained Earnings $       239,000 =71000+168000
Total Liabilities & Stockholder's Equity $       658,800